OpenStudy (anonymous):

Solve the equation. x^2 + 18x + 81 = 25

OpenStudy (ranga):

Bring the 25 over to the left side, simplify. Try to factor the quadratic expression. If that is not possible use the quadratic formula to solve for x.

OpenStudy (anonymous):

I think it's 14, -14

OpenStudy (anonymous):

use quadratic formula for x^2+18x=-56

OpenStudy (ranga):

x^2 + 18x + 81 = 25 x^2 + 18x + 81 -25 = 0 x^2 + 18x + 56 = 0 x^2 + 4x + 14x + 56 = 0 x(x + 4) + 14(x + 4) = 0 (x + 4)(x + 14) = 0 x = -4 or x = -14.

OpenStudy (anonymous):

but that gives you -187